I need your help in figuring out this problem.
X + 2/3 = 3/5

Answer by sdmmadam@yahoo.com(530)

x + 2/3 = 3/5 ----(1)
Multiply through out by 15 (which is the lcm of the drs)
15x + 2X(5) = 3X(3)
15x+10 = 9
15x = 9-10 (grouping like terms,changing sign while changing side)
15x = -1
x =-1/15 (dividing both the sides by 15)
Answer: x = (-1/15)
Verification x = (-1/15) in (1) gives
LHS = (-1/15)+2/3
=(-1+10)/15 (taking the lcm)
=9/15
=(3X3)/(3X5)
=3/5 (cancelling 3)
=RHS